A bery weak acid that is 0.424 M in concentration has a pH of 3.21. What is the Kb of the acid?
Answer #1

Concentration (C) = 0.uzum pH of wcak-acid = 32) - PH = 3 [ pke - logo (o] _3021 = [pl - loga (ouary ] pra = -loge (ke) 1 Goo